OpenStudy (anonymous):

if cathy sells 2 apple pies and 3 tacos for $7 and she sells 3 apple pies and 2 tacos for $8, what is the cost, in dollars, of three apple pies and three tacos?

hero (hero):

2a + 3t = 7 3a + 2t = 8 Solve for a and t

hero (hero):

After solving for a and t, plug the values into: 3a + 3t =